OpenStudy (anonymous):

It's not a very difficult question, but if I could have the formulas for these two that would be great, because I'm REALLY confused and I can't find them, thank you! c: Point-slope form and Slope-intercept form!

OpenStudy (anonymous):

Point-slope form: \((y-y_1)=m(x-x_1)\) Slope-intercept form: \(y=mx+b\)

OpenStudy (anonymous):

isn't standard for the same as slope-intercept then?

OpenStudy (anonymous):

O.o? yeah kinda, but if you are given a point and slope, use point-slope but if you are give the y-int and slope, use slop-int i think it's just something like a process on how you would solve it

OpenStudy (anonymous):

oh! alright, well thank you very much, a medal for you! <3

OpenStudy (anonymous):

=) np